OpenStudy (anonymous):

Trevor split his savings of $59.75 equally to buy birthday presents for each of his friends with a birthday in March. If he spent $11.95 on each present, which equation gives the number of Trevor's friends who have a birthday in March?

OpenStudy (anonymous):

Let x be the number of friends: Then 59.75/x =11.95 => 11.95x = 59.75 => 11.95x - 59.75 =0 or x = 59.75/11.95 =5 Therfore he has 5 friends.